OpenStudy (anonymous):

A particle moves along the curve of y^3=2x+1 so that it's distance from the x-axis is increasing at a constant rate of 2 units/sec. When t=0, the particle is at (0,1). (a) Find a pair of parametric functions x=x(t) and y=y(t) that describe the motion of the particle for nonnegative t. (b) Find a , the the magnitude of the particle's acceleration, when t=1.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Destinymasha

jimthompson5910 (jim_thompson5910):

Normally you let x = t, but because of that cube on the y, it'll probably be easier to make y = t. So if y = t, then y^3 = t^3 and this is equal to 2x+1. Solve t^3 = 2x+1 for x to get 2x+1 = t^3 x = (t^3 - 1)/2

jimthompson5910 (jim_thompson5910):

so x = (t^3 - 1)/2 y = t

OpenStudy (anonymous):

So that's all you have to do to find the parametric equations?

jimthompson5910 (jim_thompson5910):

yeah I think so

jimthompson5910 (jim_thompson5910):

there may be another way to do it though

OpenStudy (anonymous):

How about the second part?

jimthompson5910 (jim_thompson5910):

wait

jimthompson5910 (jim_thompson5910):

I messed up it says "When t=0, the particle is at (0,1). " but if you plug t = 0 into y = t, you don't get y = 1

OpenStudy (anonymous):

oh that's right

jimthompson5910 (jim_thompson5910):

I'm not sure how to fix it, but I'll think of a way. One sec

jimthompson5910 (jim_thompson5910):

ok fixed it

jimthompson5910 (jim_thompson5910):

what you have to do is replace every t with t+1 to make sure we're at (0,1) when t = 0

jimthompson5910 (jim_thompson5910):

so we have x = ( (t+1)^3 - 1 )/2 y = t+1

OpenStudy (anonymous):

ah i see

jimthompson5910 (jim_thompson5910):

to find the acceleration, you need to compute x'' and y'' and plug in t = 1

OpenStudy (anonymous):

so the parametric functions you just derived represented a position vector?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

ok thanks a lot

jimthompson5910 (jim_thompson5910):

np